# In Praise of Stubbornness: The Case for Cognitive-Dissonance-Aware Knowledge Updates in LLMs

### v1

@@ Abstract
Despite remarkable capabilities, large language models (LLMs) struggle to continually update their knowledge without catastrophic forgetting. In contrast, humans effortlessly integrate new information, detect conflicts with existing beliefs, and selectively update their mental models. This paper introduces a cognitive-inspired investigation paradigm to study continual knowledge updating in LLMs. We implement two key components inspired by human cognition: (1) Dissonance and Familiarity Awareness, analyzing model behavior to classify information as novel, familiar, or dissonant; and (2) Targeted Network Updates, which track neural activity to identify frequently used (stubborn) and rarely used (plastic) neurons. Through carefully designed experiments in controlled settings, we uncover a number of empirical findings demonstrating the potential of this approach. First, dissonance detection is feasible using simple activation and gradient features, suggesting potential for cognitive-inspired training. Second, we find that non-dissonant updates largely preserve prior knowledge regardless of targeting strategy, revealing inherent robustness in LLM knowledge integration. Most critically, we discover that dissonant updates prove catastrophically destructive to the model’s knowledge base, indiscriminately affecting even information unrelated to the current updates. This suggests fundamental limitations in how neural networks handle contradictions and motivates the need for new approaches to knowledge updating that better mirror human cognitive mechanisms.
Machine Learning, ICML
@@ 1 Introduction
Humans effortlessly update their knowledge as they experience the world. They seamlessly integrate new information, ignore redundant stimuli, and actively resolve conflicts with existing beliefs before updating their mental models. This cognitive flexibility stems from several key abilities. Humans exhibit (1) selective attention, focusing on novel or relevant information while filtering out irrelevant or familiar stimuli (Posner et al., 1990; Petersen & Posner, 2012; Desimone et al., 1995; Ranganath & Rainer, 2003). They readily (2) detect conflicts (Croyle & Cooper, 1983) between new information and existing knowledge and actively engage in resolving them, a process known in psychology as cognitive-dissonance (Festinger, 1957; Van Veen et al., 2009). Moreover, their brains exhibit a form of (3) adaptive plasticity, allowing for updates to neural networks that can incorporate new information while often preserving existing knowledge. While the exact mechanisms are still being investigated, this process seems to balance the stability of well-established knowledge with flexibility in the face of new or uncertain information (McClelland et al., 1995; Behrens et al., 2007).
Table 1: Taxonomy of Incremental Learning Approaches. See Appendix.A for an extended version
@@
iCaRL (Rebuffi et al., 2017) Class-incremental Replay Task-Agnostic Fixed Fixed No Rehearsal
@@
EWC (Kirkpatrick et al., 2017b) Task-incremental None Task-Aware Selective Fixed No Regularization
@@
Progressive Nets (Rusu et al., 2016) Task-incremental None Task-Aware Fixed Expanding No New Subnetworks
@@
DEN (Yoon et al., 2017) Task-incremental None Task-Aware Selective Expanding No Selective Expansion
@@
GEM (Lopez-Paz & Ranzato, 2017) Task-incremental Replay Task-Aware Constrained Fixed No Constrained Optimization
ROME (De Cao et al., 2021) Fact-incremental None Fact-Aware Localized Fixed No Rank-One Update
@@
OWM (Zeng et al., 2019) Task-incremental None Task-Aware Orthogonal Fixed No Orthogonal Projection
@@
PackNet (Mallya & Lazebnik, 2018) Task-incremental None Task-Aware Selective Fixed No Weight Masking
@@
HAT (Serra et al., 2018) Task-incremental None Task-Aware Selective Fixed No Attention Masking
@@
This paper Fact-incremental None Conflict-Aware Selective Fixed Yes Neuron-Specific Update
Despite demonstrating remarkable capabilities, Large Language Models (LLMs) are still far from such learning abilities. Current LLMs face significant challenges in real-world deployment and long-term utility due to their static nature and training paradigms. They suffer from catastrophic forgetting (Kirkpatrick et al., 2017a; Kemker et al., 2018; Li et al., 2022; Luo et al., 2024; Kotha et al., 2024), where incorporating new information often leads to the erasure of previously learned knowledge. Furthermore, LLMs engage during training in indiscriminate learning, passively accepting all training data, even when it contradicts what they already learned. Despite emergent sparsity (Jaiswal et al., 2023; Mirzadeh et al., 2024), knowledge in LLMs follows backpropagation and the objective function, with no explicit mechanism for targeted knowledge storage or retrieval. This results in a situation where all weights are potential candidates for storing knowledge, necessitating comprehensive retraining to properly incorporate new information.
In this work, we embark on a systematic empirical investigation of how LLMs handle knowledge updates, drawing inspiration from human cognitive traits. Through carefully controlled experiments, we examine (1) the feasibility of Dissonance Awareness, i.e. whether it is possible to correctly classify facts into novel, familiar, and dissonant using features extracted from the LLM. We also investigate the benefits of (2) Adaptive Plasticity by studying how different neuron targeting strategies affect knowledge retention and update. For this, we develop a simple method for tracking historical neuron usage to identify ”plastic” (rarely used) and ”stubborn” (previously used) neurons, allowing us to study how knowledge updates affect different regions of the model’s parameter space. This experimental framework lets us systematically investigate fundamental properties of knowledge integration in LLMs.
Our investigation reveals a fundamental distinction in how LLMs handle knowledge updates: the case of non-dissonant updates (adding entirely new knowledge) versus dissonant updates (modifying existing associations). While prior work has focused mostly on editing individual factual associations (Meng et al., 2022a; Mitchell et al., 2022; Meng et al., 2022c) or preserving knowledge across distinct tasks as in continual learning  (Rebuffi et al., 2017; Kirkpatrick et al., 2017b; Mallya & Lazebnik, 2018), our controlled experiments take a different approach. We systematically study how the placement of new knowledge in the network’s parameter space affects both the integration of that knowledge and its impact on existing, unrelated knowledge. As shown in Tab. 1, this positions our work uniquely: rather than proposing new editing or continual learning methods, we reveal fundamental properties about how LLMs handle knowledge integration in both dissonant and non-dissonant scenarios. Critically, our experimental design allows us to precisely track the impact of updates on a controlled set of initial knowledge, providing clear visibility into how different update strategies affect unrelated information.
Key takeaways. This leads us to uncover several fundamental properties of LLM knowledge updating: (i) dissonance awareness is feasible using simple model features, suggesting potential for cognitive-inspired training; (ii) LLMs show inherent robustness when incorporating non-dissonant information, largely preserving prior knowledge regardless of targeting strategy; (iii) avoiding heavily-used (stubborn) neurons during updates further improves this robustness, motivating adaptive plasticity in this scenario; (iv) regions of the network heavily used during pre-training are particularly effective at incorporating new knowledge, extending lottery ticket hypothesis findings (Frankle & Carbin, 2019) to language models; and most critically, (v) dissonant updates prove catastrophically destructive to unrelated knowledge, suggesting fundamental limitations in how neural networks handle contradictions: while some of our targeted update strategies show comparable performance to existing editing methods like ROME and MEMIT, all approaches fundamentally struggle with dissonant updates, suggesting the need for fundamentally different mechanisms.
@@
Implications. These findings point to concrete opportunities such as the feasibility of dissonance awareness, and the benefits of adaptive plasticity in case of non-dissonant updates. But they also reveal fundamental challenges when handling contradictory information. Current approaches essentially attempt to erase and replace old knowledge - a process we show leads to catastrophic forgetting of even unrelated information. But this contrasts sharply with human cognition, where we maintain both old and new knowledge with appropriate temporal context. Consider how humans handled learning that Pluto was no longer classified as a planet: rather than erasing our previous understanding, we maintained both pieces of knowledge, understanding their historical context and why the classification changed. Our experiments motivate the exploration of future fundamentally different mechanisms for handling contradictions - ones that can maintain and contextualize conflicting information rather than attempting to overwrite it. [^1] Code available at https://github.com/bendiogene/ConflictAwareLLM/
@@
Figure 1: Overview of our emprical investigation pipeline (i) Classification pipeline to identify novel, familiar, and dissonant information using inner or output model features. (ii) Targeted update strategies for non-dissonant updates and (iii) dissonant updates.
2 Dissonance-aware Targeted Updates
Our cognitively-inspired investigation begins with a striking empirical discovery. Fig. 3 reveals a fundamental distinction in how language models handle different types of updates: while non-dissonant information (green) can often be incorporated while preserving existing knowledge, dissonant updates (red) prove catastrophically destructive across all model scales and training approaches. This pattern emerged from our controlled experimental pipeline (Fig. 1): when training models on 2,000 initial facts followed by 1,000 new facts, we observe dramatic differences between non-dissonant and dissonant updates in both their ability to learn new information (y-axis) and preserve original knowledge (x-axis). The stark contrast between these two types of updates motivates our systematic methodology: if dissonant updates are inherently destructive, can we develop mechanisms to identify them before they occur? And can targeted plasticity strategies help protect existing knowledge? To answer these questions, we first develop methods to extract and track neural activity patterns (Sec. 2.1), which serve two purposes: enabling classification of incoming information as novel, familiar, or dissonant (Sec. 3.2), and informing targeted update strategies that carefully control where new knowledge is stored in the network (Sec. 2.3).

### v2

# In Praise of Stubbornness: An Empirical Case for Cognitive-Dissonance Aware Continual Update of Knowledge in LLMs

@@ Abstract
Through systematic empirical investigation, we uncover a fundamental and concerning property of Large Language Models: while they can safely learn facts that don’t contradict their knowledge, attempting to update facts with contradictory information triggers catastrophic corruption of unrelated knowledge. Unlike humans, who naturally resist contradictory information, these models indiscriminately accept contradictions, leading to devastating interference, destroying up to 80% of unrelated knowledge even when learning as few as 10-100 contradicting facts. To understand whether this interference could be mitigated through selective plasticity, we experiment with targeted network updates, distinguishing between previously used (stubborn) and rarely used (plastic) neurons. We uncover another asymmetry: while sparing frequently-used neurons significantly improves retention of existing knowledge for non-contradictory updates (98% vs 93% with standard updates), contradictory updates trigger catastrophic interference regardless of targeting strategy. This effect which persists across tested model scales (GPT-2 to GPT-J-6B), suggests a fundamental limitation in how neural networks handle contradictions. Finally, we demonstrate that contradictory information can be reliably detected (95%+ accuracy) using simple model features, offering a potential protective mechanism. These findings motivate new architectures that can, like humans, naturally resist contradictions rather than allowing destructive overwrites.
@@ 1 Introduction
Humans protect their knowledge by detecting and resisting contradictions. When a new statement conflicts with what we already believe, we experience a psychological discomfort known as cognitive dissonance: the uncomfortable state of holding two conflicting beliefs simultaneously (Festinger, 1957; Croyle and Cooper, 1983). This discomfort drives us to hesitate, seek additional evidence and resolve the conflict through critical evaluation rather than passive acceptance. As a result, we often maintain both versions with appropriate episodic context (“Pluto was once classified a planet; today it is not”) rather than simply overwriting existing knowledge (Van Veen et al., 2009). This cognitive “stubbornness” serves as a protective mechanism for knowledge integrity.
Current Large Language Models (LLMs) have no such contradiction filter. During gradient-based training, every sample, whether consistent or contradictory, updates the same weight space indiscriminately. This difference between human and artificial cognition led us to investigate a crucial question: what happens when LLMs encounter contradictory information? Through carefully controlled experiments across different model scales (from GPT-2 to GPT-J-6B), we systematically compare how LLMs handle two types of knowledge updates: adding entirely new facts versus updating existing knowledge with contradictory information (e.g., training that “Paris is the capital of Italy" when the model knows “Paris is the capital of France”). Our investigation reveals, for the first time, a striking and concerning property: while LLMs tend to safely learn new, non-contradictory information, attempting to update learned facts with contradictory information triggers catastrophic corruption of completely unrelated knowledge (Sec. 3). Even minimal contradictory updates (as few as 10-100 facts) can destroy up to 80% of a model’s unrelated knowledge ( Fig. 3). Importantly, this effect persists across model scales and training approaches, suggesting a fundamental limitation in how neural networks react to contradictions.
Next, the brain offers a second clue: it balances rigid and still-malleable circuits. For example, classic critical-period experiments showed that, once ocular-dominance columns are set in primary visual cortex, additional learning is minimal (Wiesel and Hubel, 1963). An analogous “use-it-early or lose-it” window shapes auditory-cortex tonotopy (Kral, 2013). Furthermore, the dentate gyrus keeps adding young plastic granule cells that excel at novel pattern separation, whereas older, less-plastic granules specialize for pattern completion, i.e., the recall of established memories (Clelland et al., 2009; Nakashiba et al., 2012). This coexistence of frozen and adaptable subnetworks points to selective plasticity as another potentially protective mechanism in biological systems, complementing cognitive dissonance. This leads us to investigate a second question: could selective plasticity help artificial systems maintain knowledge integrity? And might the impact of such targeted updates differ also between contradictory and non-contradictory information?
To investigate, we implemented an analogue of selective plasticity in LLMs by identifying historically over-used neurons, encoding old knowledge (which we make become "stubborn" in future updates) versus rarely used ones (which we leave "plastic" for future training). This led to our second key discovery: an intriguing asymmetry in how selective plasticity behaves with different types of updates (Sec. 4). While selective avoidance of stubborn neurons significantly improves old knowledge retention when adding new information (98% vs 93% with standard updates), contradictory updates trigger catastrophic interference regardless of targeting strategy. This asymmetry reveals that while selective plasticity successfully protects knowledge during non-contradictory updates, it fails when applied to contradictory overwrites, an operation that biological systems may avoid altogether. This finding points to a fundamental question: should we be trying to erase and replace conflicting knowledge, or preserve both versions with appropriate episodic context, ultimately turning all updates into non-conflicting ones? (see Implications)
Despite these catastrophic effects, conflicts remain a critical blind spot across LLM research (Tab. 1), but also deployment. Current training and post-training procedures likely suffer from this interference unknowingly, as they lack any mechanism to identify contradictory information before it causes damage. This motivated our third empirical investigation: can we at least detect when information might be contradictory before training? Through controlled experiments, we demonstrate that simple classifiers using (either internal or output) model features achieve 95%+ accuracy in distinguishing novel, familiar, and contradictory facts (Sec. 5, App.C.5). This finding reveals that LLMs encode detectable signatures of conflict even if they cannot naturally resist it, offering hope for protective mechanisms.
Table 1: Conflict awareness gap. While continual learning explores many dimensions (memory usage, selective plasticity, etc.), conflict detection remains universally absent. Model editing focuses exclusively on contradictory updates by design, making it orthogonal to our challenge of continuous mixed-content integration that deployed LLMs require. See Appendix.A for an extended version.
Examples Incremental Type Memory Usage Task Awareness Weight Plasticity Architecture Update Mechanism Conflict Detection
@@
iCaRL (Rebuffi et al., 2017) Class-incremental Replay Task-Agnostic Fixed Fixed Rehearsal No
@@
EWC (Kirkpatrick et al., 2017) Task-incremental None Task-Aware Selective Fixed Regularization No
@@
Progressive Nets (Rusu et al., 2016) Task-incremental None Task-Aware Fixed Expanding New Subnetworks No
@@
DEN (Yoon et al., 2017) Task-incremental None Task-Aware Selective Expanding Selective Expansion No
@@
GEM (Lopez-Paz and Ranzato, 2017) Task-incremental Replay Task-Aware Constrained Fixed Constrained Optimization No
@@
OWM (Zeng et al., 2019) Task-incremental None Task-Aware Orthogonal Fixed Orthogonal Projection No
@@
PackNet (Mallya and Lazebnik, 2018) Task-incremental None Task-Aware Selective Fixed Weight Masking No
@@
HAT (Serra et al., 2018) Task-incremental None Task-Aware Selective Fixed Attention Masking No
@@
ROME (De Cao et al., 2021) Fact-incremental None Fact-Aware Localized Fixed Rank-One Update N/A
@@
Implications and future work Summarizing, our human-inspired empirical findings uncover for the first time a stark difference between non-dissonant updates (which LLMs handle robustly) and dissonant updates (which trigger catastrophic corruption of unrelated knowledge). They also point to concrete opportunities such as the feasibility of dissonance detection and the benefits of selective plasticity for non-contradictory updates. Most critically, they show that the most straightforward historical approach (simply erasing and replacing old knowledge) leads to catastrophic forgetting of unrelated information. This contrasts sharply with human cognition, where we maintain both old and new knowledge with appropriate episodic (temporal) context. Consider how humans handled learning that Pluto was no longer classified as a planet: rather than completely erasing our previous understanding, we maintained both pieces of knowledge, understanding their historical context (“Pluto was once classified as a planet; today it is not”).
Our discoveries inspire an intriguing hypothesis which deserves exploration in future work: perhaps the brain’s remarkable stability emerges from three complementary mechanisms: (i) cognitive dissonance for detecting and resolving contradictions, (ii) append-only updates that preserve both old and new knowledge with appropriate episodic context (avoiding direct overwrites), and (iii) selective plasticity for protected storage. While the exact nature of biological knowledge integration remains an active area of research, our empirical findings motivate exploring fundamentally different artificial architectures—ones that might incorporate the above protective mechanisms rather than attempting to blindly overwrite existing knowledge. [^1] Code available at https://github.com/bendiogene/ConflictAwareLLM/
2 Experimental Pipeline
@@
Fig. 1 illustrates our experimental pipeline to perform (i) non-dissonant and (ii) dissonant updates (with or without selective plasticity), as well as our (iii) Classification pipeline to identify novel, familiar, and dissonant information using inner or output model features.
Figure 1: Overview of our empirical investigation pipeline
2.1 Using COUNTERFACT Dataset to Create Dissonant and Non-dissonant Updates
We design our controlled experiments using the COUNTERFACT dataset (Meng et al., 2022b), which contains approximately 17,000 factual statements along with their corresponding counterfacts, enabling controlled investigation of both non-contradictory knowledge integration and contradictory updates.
Definition of Update Types.  We distinguish between:
$\bullet$ Non-dissonant updates: Adding entirely new facts that do not contradict existing knowledge (e.g., learning “Madrid is the capital of Spain" when the model does not know it).
$\bullet$ Dissonant updates: Attempting to modify existing knowledge with contradictory information (e.g., training that “Paris is the capital of Italy" when the model knows “Paris is the capital of France").
Experimental Protocol.  Our investigation starts by establishing a baseline.
Old “unrelated” knowledge: we train models on 2,000 initial facts from COUNTERFACT until reaching near 100% accuracy. From this common baseline, we investigate two distinct update scenarios:
Non-dissonant Updates: We introduce 1,000 new facts (different general knowledge from COUNTERFACT) and train until convergence. Note that these facts are non-dissonant to both the model’s pre-training knowledge and recently learned facts. We verify learning of new facts and measure retention of old baseline facts, revealing the model’s capability for safe knowledge integration.
Dissonant Updates: After first learning 1,000 new facts (as above), we extract the subset of original facts still remembered accurately. We then introduce counterfacts that directly contradict the recently learned 1,000 facts (we also conduct experiments with smaller contradiction sets of 10 and 100). These counterfacts conflict with both the model’s pre-training knowledge and recently learned facts. This “doubly dissonant” nature arises because COUNTERFACT’s facts are general knowledge that pre-trained models were most likely exposed to during training. For instance, despite its small size, GPT-2-small already knows approximately 600 of them before any fine-tuning. We measure both learning of these new contradictory statements and retention of the remaining unrelated facts identified earlier. Finally, control experiments with a third round of non-dissonant updates, instead of dissonant, confirm the catastrophic effect is due to contradictions, not repeated updates.
Evaluation Protocol.  For all stages, unless explicitly stated otherwise, we intentionally select hyperparameters (an example in App. D.2) that ensure successful learning of the target facts. This controlled setting allows us to isolate our key question: when models successfully learn new information, how much damage occurs to unrelated knowledge? Knowledge retention is measured at the end of each training stage after convergence.
Cross-validation.  We employ 5-fold cross-validation by creating different splits of the 2,000 initial and 1,000 new facts from COUNTERFACT’s 17K facts. This ensures our findings are robust across different subsets of general knowledge facts rather than specific to particular fact selections. Notably, we do not need to explicitly control for facts being “unknown" at pre-training, as COUNTERFACT’s general knowledge nature means models have likely seen similar information during pre-training—our training simply helps them achieve reliable accuracy on these facts, and ease our measurements.
Models and Implementation.  We employ models from the GPT family to enable comparison across scales, focusing on GPT-2-small and GPT-2-xl, with additional validation on GPT-J-6B. All experiments use 5-fold cross-validation, varying the specific sets of old and new facts. Implementation uses Hugging Face Transformers on NVIDIA GPUs.
Importantly, the small dataset size (17,000 facts) creates different visibility into catastrophic forgetting across model scales. Effects are most clearly observed with GPT-2-small, where our 2000 baseline facts at each fold represent a larger portion of the model’s knowledge. In GPT-2-xl, the same number of tracked facts represents a smaller fraction of total knowledge, leading to less visible forgetting in non-dissonant cases. We therefore focus our main text on GPT-2-small results while providing aligned GPT-2-xl findings in the Appendix.
Evaluation Metrics.  We measure performance using the standard factual accuracy metric from the model editing literature: the percentage of facts correctly recalled. This allows us to track both new knowledge acquisition and old knowledge retention. As revealed in Fig. 3, plotting these two dimensions against each other exposes the stark asymmetry between dissonant and non-dissonant updates, our first key finding.
2.2 Experimenting with selective plasticity
To implement selective plasticity, we follow the experimental pipeline illustrated in Fig. 1, systematically exploring where to selectively place new knowledge based on historical neuron usage.
2.2.1 Historical tracking during baseline knowledge learning
During the initial training on the 2,000 baseline “unrelated knowledge" facts, we maintain an aggregate profile of neuronal activity by accumulating activations and gradients for each neuron at every training step. Specifically, for each neuron $n$ in the Transformer blocks (including feed-forward (MLP) layers and attention projections (Key, Query, Value matrices)), we compute $H\hat{G}_{n}$, the cumulative historical gradient magnitude over time, and $H\hat{A}_{n}$, the cumulative historical activation magnitude over time. To mitigate scale differences across layers, we also experiment with layer-wise normalization of activations and gradients before accumulation. Precise notation and more details are in Appendix B. We use the gradient [^2] We use the activations though as input features to assess feasibility of dissonance detection. historical activity to classify neurons as plastic” or stubborn” based on their past usage during baseline learning, as follows, and as visually illustrated in Fig.2.
Plastic Neurons. Neurons underutilized during baseline learning. To identify them, we rank neurons by increasing historical gradient values and select the top $N$ neurons with the lowest cumulative gradients:
$\mathcal{N}_{\text{plastic}}=\{n\mid\text{rank}(H\hat{G}_{n})\leq N\},$
where $H\hat{G}_{n}$ is the historical gradient for neuron $n$. This allows to assess whether targeting underutilized neurons can integrate new knowledge without interfering with baseline knowledge.
Stubborn Neurons. Neurons that accumulated high historical gradients during baseline learning, indicating significant involvement in storing the unrelated knowledge. We rank neurons by decreasing historical gradient values and select the top $N$ neurons:
where $|\mathcal{N}|$ is the total number of neurons, and $H\hat{G}_{n}$ is the historical gradient for neuron $n$. Updating stubborn neurons allows us to test whether overwriting neurons crucial for baseline knowledge affects new knowledge integration and baseline retention.
2.2.2 Candidate selection for new knowledge placement
Before training on the new facts (whether dissonant or non-dissonant), we identify which neurons the model naturally prefers for storing this new information. To identify them, we perform a single back-propagation pass on the new input data, without updating the model weights. We then rank neurons based on the magnitude of these gradients and select candidates accordingly.
Candidate Neurons. We rank neurons based on their gradient magnitude for the new facts and select the top $N$:
where $G_{n}^{\text{new}}$ is the gradient for neuron $n$ obtained from the back-propagation pass on the new facts. Targeting candidate neurons focuses updates on areas of the network where the model naturally wants to store the new information.
Specific Neurons. We identify neurons that the model prefers for new information (candidate) while avoiding those crucial for baseline knowledge (stubborn). For this, we first: (1) identify stubborn neurons $\mathcal{N}_{\text{stubborn}}$, using $N$ as defined earlier; (2) rank all neurons based on their gradients $G_{n}^{\text{new}}$ for the new facts; (3) select the top $N$ neurons that are not in $\mathcal{N}_{\text{stubborn}}$:
where $\mathcal{N}{\text{all}}$ is the set of all neurons ranked by their gradient magnitudes for new facts. This approach attempts to optimally place new knowledge where the model prefers while protecting baseline knowledge.
2.2.3 Targeted training and design space exploration
During training on the new facts, we perform standard forward and backward passes to compute the loss and gradients. Before the optimizer step, we modify the gradients to freeze certain neurons. Specifically, given the gradients for all parameters of a given layer, we zero-out those that do not belong to the selected set of neuron and corresponding weights. This process effectively freezes the weights of non-selected neurons, allowing for targeted updates to specific parts of the model.
We further vary the number of selected neurons to control how new information is integrated into the model while managing its impact on baseline knowledge. Our four targeting strategies systematically explore the complete design space defined by two key decisions: (1) whether to respect neurons important for old baseline knowledge, and (2) whether to follow the “model’s preferences” for storing new information. Fig. 2 illustrates the conceptual relationship between strategies within the model’s parameter space. Not shown, as a control, we add a fifth strategy which is random placement.
[…]
Strategy Avoid stubborn Follow gradient preference
Plastic
Stubborn
Candidate
Specific
Figure 2: Illustration of selective plasticity strategies based on past gradient distribution analysis (here, cumulative gradients during a GPT-2-xl training) and current preferred gradient pathways
We compare against full fine-tuning (updating all neurons) and LoRA adaptation, providing baselines for our targeted approaches. We also vary the number of updated neurons (2k-20k) to understand the trade-offs between preservation and learning capacity.
2.3 Feasibility of dissonance detection
Our third investigation examines whether LLMs encode detectable signatures of contradictory information. We frame this as a three-class classification problem: given an input sequence X, can we reliably distinguish between Novel (previously unseen), Known (familiar), and Dissonant (contradictory) information? The right side of Fig. 1 illustrates our classification pipeline.
Dataset Construction. To address the lack of truly novel facts in COUNTERFACT, we generate additional data using GPT4, transforming existing statements into plausible yet fictitious information while maintaining relational similarity. For example, "Danielle Darrieux’s mother tongue is French" becomes "Sylvan Myrthil’s mother tongue is Sylvan" (see Appendix C.1). This ensures classification cannot rely on surface features, as all three fact types share similar relational structures.
Investigation Scenarios. We investigated dissonance detection in two distinct scenarios:
$\bullet$ Pre-trained Model: Starting with GPT-2-small’s pre-existing knowledge, we identified 600 facts the model already knew reliably. From these, we constructed a balanced dataset containing these 600 known facts, their corresponding 600 contradictory facts, and 600 novel generated facts.
$\bullet$ Fine-tuned Model: We first fine-tuned GPT-2-small on 1,000 facts to establish a controlled set of known information. We then constructed a balanced dataset of 1,000 facts per class.
For each scenario, we employed 5-fold cross-validation, ensuring by design that facts used for testing share no content overlap with those used for training. Features were extracted from the corresponding model (pre-trained or fine-tuned) for each scenario.
Feature Extraction. We explored two potential sources of signatures:
$\bullet$ Internal Model Features: For each input X, we collect activations (forward pass) and gradients (backward pass, without weight updates). Following our goal of investigating feasibility, we extract simple statistical features (mean, standard deviation, min/max, quartiles) per layer, optionally normalized by historical statistics from Sec. 2.2.1. We perform ablation studies to assess the importance of different features and employ feature importance analyses to understand which aspects contribute most to the classifier’s performance.
$\bullet$ Output Distribution Features: Not shown in the main paper for lack of space, we also investigate whether model output probabilities alone could signal contradictions (reported in Appendix C.5).
For both scenarios, we employed simple classifiers (Random Forests and SVMs), optimizing hyperparameters using Bayesian search.
3 A tale of two updates: one safe, the other catastrophic
Our cognitively-inspired investigation begins with a striking empirical discovery. Fig. 3 focuses on classic and LoRa finetuning, opposing the impacts of learning dissonant (red) vs. non-dissonant (green) facts, on completely unrelated prior knowledge. While non-dissonant information (green) can often be incorporated while preserving existing knowledge, dissonant updates (red) prove catastrophically destructive across all model scales and training approaches.